OpenStudy (anonymous):

can anyone solve this: [3\sqrt{2} - 5\sqrt{7} + 2\sqrt{7} - 4\sqrt{7}\]

OpenStudy (anonymous):

\[3\sqrt{2} - 5\sqrt{7} + 2\sqrt{7} - 4\sqrt{7}\]

OpenStudy (anonymous):

you mean combine like terms?

OpenStudy (anonymous):

it's called simplifying radicals

OpenStudy (anonymous):

get \[3\sqrt{2}-7\sqrt{7}\]

OpenStudy (anonymous):

\[\sqrt{2}\] and \[\sqrt{7}\] are not like terms so you cannot combine them. simplifying radicals indeed. you are combining like terms.

OpenStudy (anonymous):

idon't even knw how to solve the problem,,better yet knw what yur tlkin bout lmaoo